OpenStudy (anonymous):

Let y=sin(2x). If delta x=0.3 and x=0, use linear approximation to estimate delta y. I found delta y to be 0.6. What is the percentage error?

OpenStudy (apoorvk):

del y = f(x + del x) - f(x) = sin(0.6) - sin(0) = 0.56 - 0 = 0.56 ^^(by using a calc.) you got 0.6 by linear approximation. so percentage error = ((0.6 -0.56()/0.6)*100 = (0.04/0.6)*100 = 6.67%

OpenStudy (anonymous):

Thank you @apoorvk !

OpenStudy (apoorvk):

No worries!